# EDA-Driven Preprocessing for SAT Solving

### v1

@@ 3.2.6 Policy
We create the target function $Q_{\hat{\theta}}(s^{t},a)$ to stabilize the training process. The parameter $\hat{\theta}$ are copied from $\theta$ after $k$ steps. The parameter $\theta$ of action-value function $Q$ is updated iteratively by loss function in Eq. (5), where $\gamma$ is a hyperparameter and represents discount factor.
@@
3.3 Cost-Customized LUT Mapping
We present a cost-customized LUT mapping approach that not only hides the internal nodes during solving, but also forms an easy-to-solve CNF instance. Unlike conventional mapping tools that focus on constructing circuits with low area or delay, our approach aims to create circuits that are easier to solve. Specifically, we begin by defining a cost metric to quantify the branching complexity of SAT solving and compute the cost values for all $4$-input LUTs. Next, we modify the cost function in the existing mapping tool to enable minimizing the overall branching complexity. Lastly, we apply the cost-customized mapper to map AIG netlist to LUT netlist and convert the LUT netlist into CNF.
3.3.1 Branching Complexity of LUT
The SAT solving process involves iteratively making variable branching decisions, where the solver selects variable assignments as either true or false. We formulate the solving process within the circuit domain. Specifically, given the logic value of a certain gate, the SAT solver is required to make decisions by selecting assignments for its fanin gates. We consider two 2-fanin LUTs, denoted as $L_{1}$ (representing an AND gate) and $L_{2}$ (representing an XOR gate), and list the truth tables for these LUTs in Figure 4.
@@ 3.3.1 Branching Complexity of LUT
Input: The input CNF instance $\phi_{in}$ or Circuit instance $\mathcal{G}_{in}$
@@
1 if input instance is in CNF then
2       $\mathcal{G}^{0}=\textit{cnf2aig}(\phi_{in})$;
@@ 5       $\mathcal{G}^{0}=\textit{aigmap}(\mathcal{G}_{in})$;
/* Explore logic synthesis recipe by RL-agent (see Section 3.2) */
@@ 19       $t=t+1$;
/* Perform cost-customized LUT mapping (see Section 3.3) */
@@ 25 $\phi_{out}=\textit{lut2cnf}(\mathcal{G}_{LUT})$;
Algorithm 1 EDA-Driven preprocessing framework for SAT solving
3.3.2 Cost-Customized Mapper
@@
In contrast to [15] that minimizes the size of the post-mapping netlist, our approach prioritizes mapping the logic into cells with lower branching complexity. Our approach is motivated by the observations that instances containing a significant proportion of XOR gates often require more time to be solved [18, 13], where the the branching complexity of XOR gate ($L_{2}$ in Figure 4) is higher than other gates.
To prioritize the LUTs during mapping process, we employ a strategy whereby we enumerate all $4$-LUTs and integrate their branching complexity into the cost function. We implement a cost-customized mapper based on mockturtle [35] by modifying the area cost to reflect the branching complexity of each LUT and fixing the the delay cost as a constraint. By executing a sequence of operations aimed at minimizing the total cost, the resulting post-mapping netlists exhibit minimal overall branching complexity. To ensure the compatibility of CNF-based SAT solver, our preprocessing strategy ends with a LUT to CNF transformation [24].
As a result, we formally depict our overall EDA-driven preprocessing framework coupled with RL-based logic synthesis exploration and cost-customized LUT mapping in Algorithm 1. The framework begins by transforming the input CNF instance $\phi_{in}$ or circuit instance $\mathcal{G}_{in}$ into AIG format, using either $cnf2aig$ tool [31] or $aigmap$ technique in ABC [7]. Then, the RL agent (elaborated in Section 3.2) proceeds to select logic synthesis operations step by step through making action $a^{t}$. During this process (line 6-16), the circuit is synthesized multiple times to create a simplified circuit. Next, the framework continues to perform cost-customized LUT mapping on the AIG netlist $\mathcal{G}^{t}$, resulting in a compact LUT netlist $\mathcal{G}$. Finally, the LUT netlist is encoded into CNF as output $\phi_{out}$. Due to these simplification approaches, $\phi_{out}$ has a lower solving complexity than $\phi_{in}$ and the subsequent SAT solving procedures can be more efficiently. We will assess the effectiveness of our preprocessing by the following experiments.
4 Experiments
4.1 Experiment Settings
We collect $200$ industrial instances in AIG format into the RL training dataset. These AIG instances are derived from logic equivalence checking (LEC) problems and only contain a single primary output. The statistics of these instances are shown in Table 1, where # Clauses shows the number of clauses after transforming the AIG instances into CNFs and Time represents the solving time without any preprocessing.
We set the maximum number of steps per episode as $T=10$ and the discount factor $\gamma=0.98$. During each training episode, the agent randomly selects an instance and makes decisions aimed at transforming the instance to minimize the number of branching times in SAT solving. We conduct a total of $10,000$ episodes, i.e., the RL agent preprocesses instances for $10,000$ times during RL training. The batch size is $32$ and learning rate is $10^{-5}$.
Table 1: Statistics of training dataset
@@ 4.1 Experiment Settings
Time (s) 2.01 1.96 6.68 0.04
@@
We opt for the ABC tool [7] to synthesis circuit and build a cost-customized mapper based on mockturtle [35] to map circuit into LUT netlists. Besides, we use Kissat SAT solver [16] to produce rewards in the RL environment and get the solving results in the following experiments. The solver runtime limitation is $1,000$s. Any cases that exceed this time limit are marked with TO (Time-out) and are considered to have a solving time of $1,000$ seconds ($\mathbf{T}_{solve}=1,000$).
As shown in Table 2, we randomly collect $5$ industrial LEC cases (I1-I5) and $8$ SAT competition benchmark cases (C1-C8) as testing data. These LEC cases share the same distribution as training samples but are not seen by the RL agent during training. We conduct the ablation study with these LEC cases in the following sections. Additionally, we list the number of gates (# Gates), the number of variables (# Vars), clauses (# Clas) and solving time $\mathcal{T}_{solve}$. It should be noted that Cases C1-C8 do not have natural circuit structures.
Table 2: Characteristics of testing cases
Case # Gates # Vars # Clas $\mathcal{T}_{solve}$
I1 43,865 42,069 105,711 322.46
I2 46,867 44,949 112,954 708.97
I3 43,825 42,038 105,629 531.94
I4 38,939 37,275 93,678 289.89
I5 31,337 30,087 75,537 172.79
C1 121,718 324,889 968.73
C2 164,084 434,279 TO
C3 27,154 70,968 153.96
C4 47,894 127,757 190.79
C5 4,435 9,814 50.69
C6 5,605 12,432 TO
C7 46,347 120,843 118.47
C8 N/A 38,921 102,787 324.97
4.2 Solving Time Comparison on LEC Cases
@@
To showcase the efficiency of our preprocessing framework, we compare the solving time with another circuit-based approach [^1] We do not include comparisons with CNF-based preprocessing approaches because our framework is not mutually exclusive with them and can be used in conjunction. In our experiment, we keep the default CNF-based preprocessing in Kissat solver. [15]. To the best of our knowledge, this is the only circuit-based SAT preprocessing technique.
Table 3 shows the solving time for $5$ industrial logic equivalence checking instances I1-I5. In the Baseline setting, the instances are solved directly. We denote the agent runtime, transformation time and solving time as $\mathcal{T}_{agent}$, $\mathcal{T}_{trans}$ and $\mathcal{T}_{solve}$, respectively, measured in seconds. The overall runtime, which sums up all three components, is denoted as $\mathcal{T}_{all}$.
Table 3: Solving time comparison between Ours and [15] on LEC cases
{adjustwidth}
-10pt0cm Baseline  [15] Ours $\mathcal{T}_{solve}$ Vars # Clas $\mathcal{T}_{trans}$ $\mathcal{T}_{solve}$ $\mathcal{T}_{all}$ Red. # Vars # Clas $\mathcal{T}_{agent}$ $\mathcal{T}_{trans}$ $\mathcal{T}_{solve}$ $\mathcal{T}_{all}$ Red. I1 322.46 5,616 54,529 5.31 51.49 56.80 82.39% 3,623 29,053 9.22 5.92 3.21 18.34 94.31% I2 708.97 6,052 60,573 5.61 147.85 153.46 78.35% 4,488 38,245 10.28 6.22 2.20 18.70 97.36% I3 531.94 5,612 54,825 5.21 109.89 115.10 78.36% 4,216 34,533 9.04 5.92 1.46 16.42 96.91% I4 289.89 5,038 49,805 4.61 90.05 94.66 67.35% 3,791 31,633 7.39 5.11 1.77 14.28 95.08% I5 172.79 4,006 38,069 3.91 38.77 42.67 75.30% 2,649 20,929 5.09 4.41 0.89 10.39 93.99% Avg. 405.21 92.54 77.16% 15.63 96.14%
From Table 3, we have two observations. First, our proposed framework shows remarkable performance in reducing the solving time. For instance I2, the baseline solving time is $708.97$ seconds, which significantly decreases to only $2.20$ seconds after applying our preprocessing framework. Second, our framework achieves overall solving time reduction by $96.14\%$ on average, surpassing the performance of another circuit-based preprocessing approach [15] by $83.11\%$, which only achieves an average reduction of $77.16\%$.
4.3 Effectiveness of Logic Synthesis Agent
@@
To investigate the effectiveness of the RL agent, we conduct an ablation study by introducing another agent that randomly selects logic synthesis operations for $T=10$ times (noted as w/o RL). The original setting with RL agent is noted as w/ RL. We test these two settings using the above $5$ industrial LEC cases and summarized the results in Table 4.
Table 4: Solving time comparison between w/ and w/o RL agent
Baseline w/o RL w/ RL
$\mathcal{T}_{solve}$ Vars # Clas $\mathcal{T}_{trans}$ $\mathcal{T}_{solve}$ $\mathcal{T}_{all}$ $\mathcal{T}_{solve}$ $\mathcal{T}_{all}$
I1 322.46 6,139 50,665 3.71 46.08 49.79 3.21 18.34
I2 708.97 6,581 55,185 4.01 73.03 77.04 2.20 18.70
I3 531.94 6,145 50,845 3.71 57.70 61.41 1.46 16.42
I4 289.89 5,444 45,545 3.21 46.98 50.19 1.77 14.28
I5 172.79 4,376 36,089 2.61 28.85 31.46 0.89 10.39
Avg. 53.98 15.63
@@
According to these results, we observe that although the agent with a random policy is capable of employing logic synthesis operations to simplify the circuit, it still falls short in terms of achieving optimal solving time reduction. Specifically, the average overall runtime ($\mathcal{T}{all}$) for the w/o RL setting is $53.98$s, which is about $2.45$x longer than that for w/ RL setting ($\mathcal{T}_{all}$=$15.63$ on average).
4.4 Effectiveness of Cost-Customized Mapper
@@
To showcase the effectiveness of our proposed cost-customized mapper, we replace our cost-customized mapper with a mapper with conventional cost metrics. We refer to the new setting as Conventional Mapper and our original setting as Our Mapper. It is worth noting that both settings utilize the same RL agent with same parameters, and thus, we disregard the agent runtime ($\mathcal{T}_{agent}$) in the reported results. The remaining results, including the transformation time and solving time, are presented in Table 5.
@@
First, there is a minor difference in the transformation time between two settings, suggesting that our cost-customized mapper does not increase the LUT mapping time. Second, the instances processed by the conventional mapper require $3.07$ seconds to solve on average, which is $60.73\%$ longer than the solving time of Our Mapper. We attribute this difference to the diverged optimization objectives in two mappers. The conventional mapper focuses on minimizing area and delay, while our cost-customized mapper directly targets the optimization of solving complexity. By prioritizing solving complexity reduction as the objective, our mapper achieves superior solving times compared to the conventional mapper.
4.5 Extrapolating on Novel Distributed Problems
Table 5: Solving time comparison between different mappers
Baseline Conventional Mapper Our Mapper
$\mathcal{T}_{solve}$ Vars # Clas $\mathcal{T}_{trans}$ $\mathcal{T}_{solve}$ $\mathcal{T}_{trans}$ $\mathcal{T}_{solve}$
I1 322.46 3,160 31,281 5.62 4.43 5.92 3.21
I2 708.97 4,112 41,873 6.12 4.41 6.22 2.20
I3 531.94 3,849 37,329 5.61 2.91 5.92 1.46
I4 289.89 3,478 34,013 5.11 2.50 5.11 1.77
I5 172.79 2,311 22,473 4.31 1.10 4.41 0.89
Avg. 5.35 3.07 5.51 1.91
Table 6: Solving time comparison between Ours and [15] on SAT competition benchmarks
{adjustwidth}
-30pt0cm Case Baseline  [15] Ours $\mathcal{T}_{solve}$ Vars # Clas $\mathcal{T}_{trans}$ $\mathcal{T}_{solve}$ $\mathcal{T}_{all}$ Red. # Vars # Clas $\mathcal{T}_{agent}$ $\mathcal{T}_{trans}$ $\mathcal{T}_{solve}$ $\mathcal{T}_{all}$ Red. C1 968.73 31,343 373,249 18.03 815.72 833.76 13.93% 40,565 429,333 100.16 21.74 148.15 270.05 72.12% C2 TO 32,075 352,309 49.99 TO TO 0.00% 42,732 428,121 215.76 17.93 531.15 764.84 23.52% C3 153.96 5,891 84,689 2.71 122.20 124.91 18.87% 6,102 85,101 5.54 4.71 106.88 117.13 23.92% C4 190.79 6,146 83,589 17.13 199.03 216.16 -13.30% 9,955 123,797 10.03 6.11 136.13 152.27 20.19% C5 50.69 1,298 9,701 0.70 46.58 47.29 6.72% 1,295 9,813 2.84 1.60 31.16 35.60 29.77% C6 TO 1,629 12,293 0.50 592.06 592.56 40.74% 1,658 12,493 4.15 1.61 380.75 386.51 61.35% C7 118.47 6,107 71,285 20.84 194.05 214.89 -81.39% 10,222 118,981 15.03 6.12 18.93 40.08 66.17% C8 324.97 10,063 136,713 7.79 144.00 151.79 53.29% 9,417 114,337 14.10 5.91 25.24 45.26 86.07% Avg. 475.95 397.67 16.45% 226.47 52.42%
To further investigate the generalization ability of our proposed framework, we perform an evaluation using $8$ instances (C1-C8) collected from SAT competition benchmarks [2]. These instances are originally represented as CNF and exhibit diverse distributions.
The solving results compared with [15] are listed in Table 6. We have two observations based on this table. Firstly, our preprocessing framework demonstrates an average reduction in $\mathcal{T}_{all}$ of $52.42\%$. In contrast,  [15] only achieves an average reduction of $16.45\%$ on $\mathcal{T}_{all}$, making our framework $2.19$x more efficient in comparison. It is noteworthy that the compared approach fails to solve instance C2 within the time limit, while the same instance requires only $764.84$s to be solved after processing by our framework.
Secondly, the instances transformed by our framework may exhibit a larger number of clauses (# Clas) compared to the another preprocessing [15]. For example, after being processed by our framework, instance C1 contains $429,333$ clauses, significantly exceeding the # Clas of $373,249$ in [15]. We argue that the size of the problem does not always correlate with solving time. Despite the increase in problem size, we effectively reformulate the instance using EDA tools, resulting in a simplified CNF that is easy to solve.
To sum up, our EDA-driven SAT preprocessing can be generalized to other novel instances derived from diverse problem sets.
4.6 Discussion of CNF to Circuit Transformation
While our proposed EDA-driven preprocessing framework achieves an average reduction in solving time of $52.42\%$ for the CNF-based SAT benchmarks, there are still performance gaps compared to the LEC cases, which shows a higher solving time reduction of $96.14\%$. We attribute the observation to the absence of topological structure in the flatten CNF representations and the limitations of existing CNF-to-circuit transformation techniques.
According to Algorithm 1, if the input instance is in CNF, it is required to be transformed into circuit by recovering topological information. To the best of our knowledge, the only available tool for this transformation is cnf2aig [31] (part of AIGER [5]), which iteratively selects the maximum set of clauses to match the behavior of a logic gate. From Table. 7, we list the number of gates (# Gates), logic levels (# Levs) before preprocessing and the number of LUTs (# LUTs) after preprocessing. First, the CNF-based cases (C1-C8) lacking natural topological structures are transformed into narrow AIG circuits before preprocessing, which have thousands of logic levels but only $2.66$ gates per level. Consequently, there is a significant disparity between the transformed AIG circuits and real circuit designs, rendering the former unsuitable for conventional logic synthesis operations. Second, after the transformation from CNF to circuit, the LUT netlists of the CNF-based instances exhibit a higher degree of flattening. On average, these transformed instances have approximately $482.62$ LUTs per level, which is significantly larger compared to the circuit-based instances (I1-I5) that have an average of $79.33$ LUTs per level.
To address these challenges and bridge the performance gaps, future research will focus on improving CNF-to-circuit transformation tool. Such tool should involve more effective algorithms that can formulate more realistic circuit structures from pure CNF instances.
Table 7: Circuit size after and before preprocessing
Case Before Preprocessing After Preprocessing
# Gates # Levs Gates / Lev # LUTs # Levs LUTs / Lev
I1 43,865 194 226.11 3,623 47 77.09
I2 46,867 200 234.34 4,488 51 88.00
I3 43,825 192 228.26 4,216 48 87.83
I4 38,939 184 211.63 3,791 47 80.66
I5 31,337 168 186.53 2,649 42 63.07
Avg. 217.37 79.33
C1 181,714 48,122 3.78 40,565 17 2386.18
C2 243,360 65,742 3.70 42,732 17 2513.65
C3 36,078 17,340 2.08 6,102 12 508.50
C4 66,960 27,036 2.48 9,955 16 622.19
C5 5,642 1,978 2.85 1,295 10 129.50
C6 7,140 2,506 2.85 1,658 11 150.73
C7 63,280 27,126 2.33 10,222 13 786.31
C8 55,882 19,964 2.80 9,417 13 724.38
Avg. 2.66 482.62
5 Conclusion
This paper introduces an innovative EDA-driven preprocessing framework designed to optimize SAT instances prior to solving. Our approach begins by transforming CNF into circuit format, enabling the application of circuit optimization techniques to SAT problems. We employ advanced logic synthesis methods to reformulate the circuit, focusing on identifying the most effective strategies for minimizing solving time. Additionally, our framework customizes cost metrics in technology mapping tools and leverages LUT mapping to produce instances that are easier to solve. Seamlessly integrating as a plug-in within the existing SAT solving pipeline, our framework ensures compatibility and enhances efficiency. Experimental results demonstrate a significant reduction in solving time for both circuit-based and novel distributed CNF instances, showcasing the framework’s effectiveness and versatility.

### v2

@@ III-B6 Policy
We create the target function $Q_{\hat{\theta}}(s^{t},a)$ to stabilize the training process. The parameter $\hat{\theta}$ is copied from $\theta$ after $k$ steps. The parameter $\theta$ of action-value function $Q$ is updated iteratively by loss function in Eq. (5), where $\gamma$ is a hyperparameter and represents discount factor.
@@
III-C Cost-Customized LUT Mapping
We present a cost-customized LUT mapping approach that not only hides the internal nodes during solving, but also forms an easy-to-solve instance. Unlike conventional mapping tools that focus on constructing circuits with low area or delay, our approach aims to create circuits that are easier to solve. Specifically, we begin by defining a cost metric to quantify the branching complexity of SAT solving and compute the cost values for all $4$-input LUTs. Next, we modify the cost function in the existing mapping tool to enable minimizing the overall branching complexity. Lastly, we apply the cost-customized mapper to map AIG netlist to LUT netlist and convert the LUT netlist into CNF.
III-C1 Branching Complexity of LUT
The SAT solving process involves iteratively making variable branching decisions, where the solver selects variable assignments as either true or false. We formulate the solving process within the circuit domain. Specifically, given the logic value of a gate, the SAT solver makes decisions by selecting assignments for its fanin gates. We consider two 2-fanin LUTs, denoted as $L_{1}$ (representing an AND gate) and $L_{2}$ (representing an XOR gate), and list the truth tables for these LUTs in Fig. 3.
@@ III-C1 Branching Complexity of LUT
Input: The input circuit instance $\mathcal{G}_{in}$
@@
1 if input instance is in AIG then
2       $\mathcal{G}^{0}=\mathcal{G}_{in}$;
@@ 5       $\mathcal{G}^{0}=\textit{aigmap}(\mathcal{G}_{in})$;
/* Explore logic synthesis recipe by RL-agent (see Section III-B) */
@@ 19       $t=t+1$;
/* Perform cost-customized LUT mapping (see Section III-C) */
@@ 25 $\phi_{out}=\textit{lut2cnf}(\mathcal{G}_{LUT})$;
Algorithm 1 CSAT preprocessing framework
III-C2 Cost-Customized Mapper
@@
In contrast to [8] that minimizes the size of the post-mapping netlist, our approach prioritizes mapping the logic into cells with lower branching complexity. Our approach is motivated by the observations that instances containing a significant proportion of XOR gates often require more time to be solved [34, 35], where the branching complexity of XOR gate ($L_{2}$ in Fig. 3) is higher than other gates.
To prioritize the LUTs during mapping process, we employ a strategy to enumerate all $4$-LUTs and integrate their branching complexity into the cost function. We implement a cost-customized mapper based on mockturtle [13] by modifying the area cost to reflect the branching complexity of each LUT and fixing the delay cost as a constraint. By executing a sequence of operations aimed at minimizing the total cost, the resulting post-mapping netlists exhibit minimal overall branching complexity.Our preprocessing strategy ends with a LUT to CNF transformation [36] for compatibility.
As a result, we formally depict our overall circuit-based preprocessing framework coupled with RL-based logic synthesis exploration and cost-customized LUT mapping in Algorithm 1. The framework begins by transforming the input circuit instance $\mathcal{G}_{in}$ into AIG format, using $aigmap$ technique in ABC [37]. Then, the RL agent (elaborated in Section III-B) proceeds to select logic synthesis operations step by step through making action $a^{t}$. During this process (line 6-16), the circuit is synthesized multiple times to create a simplified circuit. Next, the framework continues to perform cost-customized LUT mapping on the AIG netlist $\mathcal{G}^{t}$, resulting in a compact LUT netlist $\mathcal{G}$. Finally, the LUT netlist is encoded into CNF as output $\phi_{out}$. Due to these simplification approaches, $\phi_{out}$ has a lower solving complexity than $\phi_{in}$ and the subsequent SAT solving procedures can be more efficient. We will assess the effectiveness of our preprocessing by the following experiments.
IV Experiments
(a) Kissat
(c) Cadical
Figure 4: Runtime Comparison with Baseline and Comp.
IV-A Experiment Settings
TABLE I: Statistics of training dataset
@@ IV-A Experiment Settings
Time (s) 2.01 1.96 0.04 6.68
We collect $200$ easy (solving time $0.04s-6.68s$) instances into the RL training dataset and $300$ hard instances for testing. All instances are in AIG format and derived from both industrial logic equivalence checking (LEC) and automatic test pattern generation (ATPG) problems. For LEC instances, we modify original industrial datapath circuits and connect their primary outputs (POs) through XOR gates to form CSAT instances, where satisfiability indicates non-equivalent functionality between circuits. For ATPG instances, we introduce stuck-at faults into industrial circuits and connect the POs of faulty and fault-free circuits through XOR gates, where satisfiable assignments serve as test patterns for fault detection. We construct $200$ LEC instances and $100$ ATPG instances in total, where 200 easy instances are selected for RL training based on their clause count after CNF transformation and solving time without preprocessing. The statistics of training instances are shown in Table I, where # Clauses shows the number of clauses after transforming AIG instances into CNFs and Time represents the solving time without any preprocessing.
In the RL setting, we assign the maximum number of steps per episode as $T=10$ and the discount factor $\gamma=0.98$. During each training episode, the agent randomly selects an instance and makes decisions aimed at transforming the instance to minimize the number of branching times in SAT solving. We conduct a total of $10,000$ episodes, i.e., the RL agent preprocesses instances for $10,000$ times during RL training. The batch size is $32$ and learning rate is $10^{-5}$.
@@
We opt for the ABC tool [37] to synthesize the circuit and build a cost-customized mapper based on mockturtle [13] to map the circuit in LUT netlists. The solver runtime limitation is $1,000$s. Any cases that exceed this time limit are marked with TO (Time-out) and are considered to have a solving time of $1,000$ seconds ($\mathcal{T}_{solve}=1,000s$). Each instance solving is conducted on a single core of Intel Xeon CPU E5-2630.
IV-B Runtime Comparison
Figure 4 illustrates the relationship between the number of instances solved and overall runtime (including RL model inference time, transformation time and solving time). The Baseline setting represents the conventional solving pipeline, encoding the circuit-based instances directly into CNFs. Our experiment evaluates the effectiveness of our proposed framework using both Kissat solver (version 4.0.0 [38]) and CaDiCaL solver (version 2.0.0 [39]), with our specific settings distinguished by the orange line (denoted as Ours).
@@ IV-B Runtime Comparison
To showcase the efficiency of our preprocessing framework, we compare the solving time with another circuit-based approach [^1] We do not include comparisons with CNF-based preprocessing approaches because our framework is not mutually exclusive with them and can be used in conjunction. Pratically, we keep the default CNF-based preprocessing. [8], denoted as Comp. . To the best of our knowledge, although such approach has been around for over a decade, this is the only circuit-based SAT preprocessing technique.
According to Figure 4, our proposed circuit-based proprocessing framework can significantly reduce runtime. To be specific, solving these 300 instances necessitates 19,422.38 seconds for the CaDiCaL Baseline pipeline and 11,073.88 seconds for the comparative pipeline (Comp.). In contrast, our method completes the task in a total of 7,179.80 seconds, marking a reduction of $63.03\%$ and $35.16\%$ in solving time compared to Baseline and Comp., respectively.
IV-C Ablation Studies
In the following experiments, we investigate the effectiveness of our RL-based agent and cost-customized mapper. We keep the same experiment environments and test cases above.
Figure 5: Ablation Study
IV-C1 Effectiveness of Logic Synthesis Agent
@@
To investigate the effectiveness of the RL agent, we conduct an ablation study by introducing another agent that randomly selects LS operations for $T=10$ times (noted as w/o RL). The original setting with RL agent is noted as Ours. We test these two settings and summarized the results in Figure 5.
@@
We observe that although the agent with a random policy is capable of employing LS operations to simplify the circuit, it still falls short in terms of achieving optimal solving time reduction. To be specific, solving these 300 instances necessitates 7329,96 seconds for the w/o RL and $6454.02$ seconds for our w/ RL agent, marking a reduction of $11.95\%$.
IV-C2 Effectiveness of Cost-Customized Mapper
@@
To showcase the effectiveness of our proposed cost-customized mapper, we replace our cost-customized mapper with a mapper with conventional cost metrics. We denote the new setting as C. Mapper and our original setting as Ours. It is worth noting that both settings utilize the same RL agent with same parameters, and thus, we disregard the agent runtime ($\mathcal{T}_{agent}$) in the reported results.
@@
As shown in Figure 5, the 300 instances processed by the conventional mapper take 9732.64 seconds to solve, which is 50.80% longer than the solving time of Our Mapper. We attribute this difference to the divergent optimization objectives: the conventional mapper minimizes area and delay, while our cost-customized mapper targets solving complexity directly.
V Conclusion and Future Work
This paper introduces a novel circuit-based preprocessing framework for circuit satisfiability problem, integrating logic synthesis with reinforcement learning to transform circuits into solver-friendly formats. Key innovations include an RL-guided logic synthesis pipeline for optimizing circuit transformations and a cost-customized LUT mapping strategy that minimizes branching complexity during SAT solving. Experiments on industrial benchmarks demonstrate significant runtime reductions—up to 63% compared to baseline approaches—while maintaining compatibility with state-of-the-art solvers. Ablation studies validate the effectiveness of both the RL agent and cost-customized mapper in improving solving efficiency.
This work highlights the potential of combining EDA innovations with machine learning to advance SAT-solving methodologies. Beyond CSAT, the proposed framework offers a promising direction for tackling other SAT variants and related combinatorial optimization problems, enabling more efficient and scalable solutions. Future work will explore extending this approach to sequential circuits and broader EDA applications.
Acknowledgments
This work was partly supported by the General Research Fund of the Hong Kong Research Grants Council (RGC) under Grant No. 14212422 and 14202824 and partly by the National Technology Innovation Center for EDA.